# MoS: Unleashing Parameter Efficiency of Low-Rank Adaptation with Mixture of Shards

### v1

@@ A.2 Hyperparameter Configurations
We utilize LLaMA2-7B and 13B (Touvron et al., 2023) as the foundation models for our experiments, which are conducted on a single NVIDIA A100-40G GPU. Each experimental configuration is repeated two times with the random seed as 0 and 1, respectively, before reporting the average performance. The detailed procedures for finetuning and evaluation are outlined below.
@@
To reduce memory cost during finetuning, we follow the approach outlined in QLoRA (Dettmers et al., 2023), loading all pretrained models in a 4-bit NormalFloat format and utilizing a Paged AdamW Optimizer. Specifically, we utilize 4-bit quantized versions of LLaMA2-7B and 13B models in our experiments. Next, we apply LoRA to all linear layers within the Transformer blocks, including the query, key, value, output, up, gate, and down projection weights, setting the scaling factor $\alpha$ to 16 and the dropout rate to 0.1. For more efficient finetuning, we also follow the configuration from Wang et al. (2024b) to set a batch size of 16 and the maximum sequence length to 512, truncating samples during preprocessing if needed. We also cap the maximum gradient norm at 0.3 to enhance training stability. Each model undergoes 10,000 steps of finetuning using a linear learning rate scheduler with a warmup ratio of 3%. Additionally, we search for the optimal learning rate for LoRA, and apply this value to both LoRA and MoS. Specifically, with a LoRA rank of 8, we search for the best learning rate from {2e-5, 5e-5, 1e-4, 2e-4, 5e-4, 1e-3, 2e-3}. Our preliminary experiments demonstrate that 2e-4 performs the best.
@@
For inference, we leverage vLLM (Kwon et al., 2023), an efficient inference and serving library for LLMs, which greatly speeds up the generation process with minor influence in performance. We also employ greedy decoding with a maximum length of 512. For the details of scoring calculations, we closely follow the GitHub repository of TULU [^4] https://github.com/allenai/open-instruct.

### v2

@@ 6 Conclusion
7 Acknowledgement
This work was supported in part by Hong Kong Innovation and Technology Support Programme Platform Research Project fund (ITS/269/22FP), RGC grants 17204423, 17205824, C7004-22G (CRF), and C5032-23G (CRF).
@@ A.2 Hyperparameter Configurations
We utilize LLaMA2-7B, 13B (Touvron et al., 2023), and LLaMA3.2-3B (Dubey et al., 2024) as the foundation models for our experiments, which are conducted on a single NVIDIA A100-40G GPU. Each experimental configuration is repeated two times with the random seed as 0 and 1, respectively, before reporting the average performance. The detailed procedures for finetuning and evaluation are outlined below.
@@
To reduce memory cost during finetuning, we follow the approach outlined in QLoRA (Dettmers et al., 2023), loading all pretrained models in a 4-bit NormalFloat format and utilizing a Paged AdamW Optimizer. Specifically, we utilize 4-bit quantized versions of LLaMA2-7B, 13B, and Llama3.2-3B models in our experiments. Next, we apply LoRA to all linear layers within the Transformer blocks, including the query, key, value, output, up, gate, and down projection weights, setting the scaling factor $\alpha$ to 16 and the dropout rate to 0.1. For more efficient finetuning, we also follow the configuration from Wang et al. (2024b) to set a batch size of 16 and the maximum sequence length to 512, truncating samples during preprocessing if needed. We also cap the maximum gradient norm at 0.3 to enhance training stability. LLaMA2-7B and 13B undergo 10,000 steps of finetuning using a linear learning rate scheduler with a warmup ratio of 3%, while LLaMA3.2-3B is finetuned for one epoch in each task. Additionally, we search for the optimal learning rate for LoRA, and apply this value to both LoRA and MoS. Specifically, with a LoRA rank of 8, we search for the best learning rate from {2e-5, 5e-5, 1e-4, 2e-4, 5e-4, 1e-3, 2e-3}. Our preliminary experiments demonstrate that 2e-4 performs the best.
@@
For inference, we leverage vLLM (Kwon et al., 2023), an efficient inference and serving library for LLMs, which significantly speeds up the generation process with minimal impact on performance. We also employ greedy decoding with a maximum length of 512. For the details of scoring calculations, we closely follow the GitHub repository of TULU [^6] https://github.com/allenai/open-instruct.
Appendix B Additional Analysis
B.1 Theoretical Motivations of Differentiation Strategies
As shown in Sec. 2, differentiation reverses the detrimental effects of the the pure sharing mechanism. Here we demonstrate the design motivation of each specific differentiation strategies under the guidance of combinational diversity. Specifically, we approximately conceptualize differentiation as the combinational diversity (i.e., number of potential combinations) of each low-rank matrix pair. Aligning with the notations in Sec. 3, let $L$ denote the number of Transformer blocks, $r$ the rank, $e$ the equivalent rank of LoRA in terms of trainable parameters, and $l$ the number of shards per vector. In the case of pure sharing, the potential combinations for each low-rank matrix pair can be expressed as $C^{Le}_{Le}=1$, since all the parameters are shared in the same way for each pair. Subset selection enhances the number of combinations to $C^{Le}_{r}$ by selecting a subset of vector pairs for each low-rank matrix pair. Pair dissociation rapidly increases the combinational diversity by separating vector pairs into distinct pools, allowing for independent selection of vectors for each matrix in a low-rank pair, resulting in a combination count of $C^{Le}_{r}\times C^{Le}_{r}$. Vector sharding, which breaks down vectors into smaller shards, further amplifies the number of combinations to $C^{Lle}_{rl}\times C^{Lle}_{rl}$, given that $C^{Le}_{r}<C^{Lle}_{rl}$ when $r<Le$ and $l>1$. Compared to LoRA, pure sharing shares all the parameters across layers, resulting in performance degradation. Shard privatization, however, partially reverses this trend by dividing the global pool into public and private sections for improved differentiation.
B.2 Further Experiments on Differentiation
Compared to the pure sharing baseline, random scaling only introduces noise to the initialization of scalers, resulting in slight performance improvement. As a more aggressive measure, subset selection randomly masks specific vector pairs, and outperforms pure sharing remarkably. To further confirm this conclusion, we also extend these experiments to the LLaMA3.2-3B model (Dubey et al., 2024). As detailed in Table 4, random scaling still exhibits slight improvement over pure sharing, while subset selection boosts the performance with a margin over 0.8% on average.
Table 4: Results of LLaMA3.2-3B with different sharing and differentiation methods across diverse instruction following datasets. “+ Random Scaling” and “+ Subset Selection” denote the individual integration of them into the “Pure Sharing” scheme, respectively.
Method Rank # Param. MMLU BBH GSM8K TyDi QA HumanEval Avg.
EM EM EM F1 EM P@1
LoRA 2 3.04M 51.97 37.07 33.74 62.59 44.39 31.17 43.49
Pure Sharing 56 3.04M 51.25 38.73 31.92 61.86 45.17 30.46 43.23
+ Random Scaling 56 3.04M 51.35 38.81 32.75 62.18 45.11 30.52 43.45
+ Subset Selection 56 3.04M 51.86 39.88 33.89 62.28 45.31 31.14 44.06
B.3 Robustness Analysis
Performance Robustness.
We further repeat the experiments with 4 random seeds (i.e., 0, 1, 2, 3) and the LLaMA3.2-3B model. As shown in Table 5, MoS with an equivalent parameter budget to LoRA with the rank of 8 performs competitively with LoRA with the rank of 64. This validates an 8×8$\times$8 × parameter reduction achieved by MoS again, aligning with our previous conclusions. Meanwhile, compared to LoRA, MoS exhibits comparable (even lower on average) standard deviations, and provides a higher average value, indicating its similar stability and better performance.
Table 5: Results of LLaMA3.2-3B across multiple instruction-following datasets using different methods. The abbreviations “# Param.” and “Avg.” refer to “Parameter Count” and “Average”, respectively, while the subscripts denote the standard deviations.
Method Rank # Param. MMLU BBH GSM8K TyDi QA HumanEval Avg.
EM EM EM F1 EM P@1
LoRA 8 12.16M $\text{52.35}_{\pm 0.25}$ $\text{38.74}_{\pm 0.89}$ $\text{37.19}_{\pm 1.59}$ $\text{63.46}_{\pm 1.06}$ $\text{46.09}_{\pm 1.14}$ $\text{30.94}_{\pm 0.21}$ $\text{44.79}_{\pm 0.86}$
64 97.26M $\text{52.36}_{\pm 0.23}$ $\text{39.70}_{\pm 0.52}$ $\text{37.98}_{\pm 0.72}$ $\text{64.04}_{\pm 1.51}$ $\text{46.53}_{\pm 1.89}$ $\text{31.83}_{\pm 0.21}$ $\text{45.41}_{\pm 0.85}$
MoS 16 12.16M $\text{52.41}_{\pm 0.17}$ $\text{40.14}_{\pm 1.01}$ $\text{37.89}_{\pm 0.71}$ $\text{63.78}_{\pm 1.05}$ $\text{46.31}_{\pm 1.21}$ $\text{31.78}_{\pm 0.23}$ $\text{45.38}_{\pm 0.73}$
Hyperparameter Robustness.
To demonstrate the robustness of hyperparameters, we perform a grid search for the configurations of private rank and shard number per vector with 4 seeds (i.e., 0, 1, 2, 3) on the LLaMA3.2-3B model and BBH benchmark. As listed in Table 6, as the number of shards increases to provide more differentiation, the optimal private rank tends to decrease. Generally, shard numbers of 4 or 8 yield the best results. From another perspective, for any given private rank, there always exists a suitable range of shard numbers that consistently produce remarkable results (i.e., $\geq$ 39.8%), demonstrating the robustness of private rank.
Significance Test.
Furthermore, we conduct the significance test between LoRA and MoS on Rank = 2 and 8 across all benchmarks in our main experiments on the LLaMA2-7B model. As shown in Table 7, the p-values between LoRA and MoS with both high and low trainable parameter budgets are smaller than 5%, indicating the statistical significance of our results.
Table 6: Results of MoS with the LLaMA3.2-3B model and BBH benchmark across different shards per vector and private ranks.
Shards per Vector Private Rank
1 3 5 7
1 38.9 39.3 39.1 39.3
2 38.6 39.3 39.7 39.5
4 39.3 39.8 40.0 39.6
8 39.6 39.8 39.6 38.8
16 39.8 39.3 39.3 38.6
Table 7: P-values of the significance test of LoRA and MoS on the LLaMA2-7B model across two different trainable parameter budgets.
Method Rank # Param. p-value
LoRA vs. MoS 2 5.00M 0.03%
LoRA vs. MoS 8 19.99M 1.29%
Appendix C Limitations
Slightly more GPU consumption is required for finetuning.
Compared to LoRA with the same trainable parameter count, MoS increases the rank several times, leading to similar GPU consumption to LoRA with the same rank. However, MoS does provide better performance. We also finetune the LLaMA3.2-3B model for one epoch across various tasks, and record the time consumption of LoRA and MoS in Table 8. It is worth noticing that MoS only incurs 2.80% more finetuning time.
The routing mechanism may incur more inference latency.
Targeting this drawback, we give up existing activation-based routing mechanisms, because the routing operation has to wait for the activations, resulting in larger latency. Instead, we intentionally adopt the index-based routing mechanism so that precomputing can be used to prepare the whole low-rank matrices in parallel to the computation of preceding transformer blocks. This could circumvent the latency of routing operation, and apply all existing inference techniques of LoRA.
Table 8: Comparison of fine-tuning time consumption (in hours) between LoRA and MoS [^8] Due to the same finetuning process as MMLU benchmark, TyDi QA is neglected here to avoid repetitive computation..
Method Rank # Param. MMLU BBH GSM8K Codex-Eval Avg.
LoRA 8 12.16M 1.50 1.47 1.82 0.21 1.25
MoS 8 12.16M 1.54 1.52 1.86 0.22 1.29